### v2

# The discrete periodic Pitman transform: invariances, braid relations, and Burke properties

Abstract.
We develop the theory of the discrete periodic Pitman transform, first introduced by Corwin, Gu, and the fifth author. We prove that, for polymers in a periodic environment, single-path and multi-path partition functions are preserved under the action of this transform on the weights in the polymer model. As a corollary, we prove that the discrete periodic Pitman transform satisfies the same braid relations that are satisfied for the full-line Pitman transform shown by Biane, Bougerol, and O’Connell. Combined with a new inhomogeneous Burke property for the periodic Pitman transform, we prove a multi-path invariance result for the periodic inverse-gamma polymer under permutations of the column parameters. In the limit to the full-line case, we obtain a multi-path extension of a recent invariance result of Bates, Emrah, Martin, Seppäläinen, and the fifth author, in both positive and zero-temperature.
1. Introduction
The study of the Pitman transform traces back to Pitman’s celebrated 2MX2M-X theorem [Pit75]. Since then, there have been numerous studies around this transform and its many variants. In the semi-discrete case, the positive-temperature lifting of the full-line Pitman transform was introduced by [MY01, O’C03] in the context of Brownian queuing process and later by [O’C12] in the setting of semi-discrete directed polymers. In the fully discrete case, the discrete geometric Pitman transform was connected to the geometric lifting of the classical Robinson-Schensted-Knuth (RSK) correspondence [Kir01] and studied by [NY04, COSZ14, OSZ14, Cor21] in the context of discrete directed polymer models.
Despite arising in different contexts, these Pitman transforms share a rich set of both algebraic and probabilistic properties that have enabled significant progress in the study of stochastic integrable systems, including directed percolation, random polymers, the KPZ equation, and so on. For example, the Burke property–the analogous product measure preservation property of Burke’s theorem for $M/M/1$ queues [Bur56]–has been proved in several contexts (see [HB76, OY01, OY02, KOR02, Mar09, DMO05, AD95] as well as [EJS24, BEM+25a] in an inhomogeneous setting) and has been used to determine the fluctuation exponents in several last-passage and polymer settings [CG06, BCS06, SV10, Sep12, EJS23, LNS23]. Moreover, the joint invariance of last-passage values/polymer partition functions under the zero-temperature/positive-temperature Pitman transforms provides key tools to the construction of central limiting objects in the KPZ universality class [DOV22] and for proving convergence of many different models to these limiting objects [DV21, Wu23, ACH24, Zha25].
The purpose of this paper is to demonstrate the rich algebraic and probabilistic structure of a periodic version of the Pitman transform, recently introduced in [CGS24]. Our first main result, Theorem 1.4, gives an invariance result for directed polymers in a periodic environment under a periodic Pitman transformation of the column weights. Previous results of this flavor in a non-periodic setting have appeared in several contexts and are discussed after the statement of the theorem. The key machinery is the encoding of geometric RSK in terms of matrix relations from [NY04]. Corwin [Cor21] previously used these tools to prove the analogous invariance for full-space polymers. The next main result, Theorem 1.6, uses this polymer invariance result to prove braid relations for the periodic Pitman transform, analogous to the relations enjoyed by the full-line Pitman transform [BBO05]. More specifically, we consider a collection of operators on the space $(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ of bi-infinite sequences of $N$-periodic vectors and operators $\mathcal{P}_{k}$, which replace the $k$th and k+1k+1st vector with their Pitman transform. The braid relations, along with the fact that each $\mathcal{P}_{k}$ is an involution (proved in [CGS24], see Lemma A.2) shows that this collection of operators generates a copy of the infinite symmetric group. We then utilize the invariance of partition functions and a new inhomogeneous Burke property (Proposition 4.1) to prove a distributional invariance of the inverse-gamma polymer partition function under the permutation of the row and column parameters in both the periodic and full-line cases (Theorems 1.7 and 1.8). In the full-line case, this gives a multi-path and positive temperature extension of a recent result of Bates, Emrah, Martin, Seppäläinen, and the fifth author [BEM+25b]. This result bears similarity to an invariance found previously by Dauvergne [Dau22]. See the discussion before and after Theorem 1.8 for more on the connection to other works. Each of our main results has an analogue in zero-temperature. These results are stated in Section 1.6.
1.1. The positive-temperature discrete periodic Pitman transform
For $N\in\mathbb{N}$, let $\mathbb{R}^{\mathbb{Z}_{N}}$ denote the set of $N$-tuples, together with periodic indexing. That is, for $\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}$, $\mathbf{X}=(X_{1},X_{2},\ldots,X_{N})$, where the indices represent elements of $\mathbb{Z}_{N}$ with addition modulo $N$ (see Figure 1). For a vector $\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}$ and indices $i,j\in\mathbb{Z}_{N}$, we set
$\mathbf{X}_{(i,j]}:=\sum_{\ell=i+1}^{j}X_{\ell},\quad\text{and}\quad\mathbf{X}_{[i,j]}:=\sum_{\ell=i}^{j}X_{\ell},$ (1.1)
where the sums are taken in cyclic order. We follow the convention that $\mathbf{X}_{(i,i]}=0$.
j][i
Figure 1. Representation of indices $[i,j]$, demonstrating periodicity.
The positive-temperature discrete periodic Pitman transform is the map $\mathcal{P}:\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}\to\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$ defined by
$\mathcal{P}(\mathbf{W}_{1},\mathbf{W}_{2})$ $=(T(\mathbf{W}_{2},\mathbf{W}_{1}),D(\mathbf{W}_{2},\mathbf{W}_{1})),\quad\text{where}$ (1.2)
$T(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{1,i-1}+\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{[i-1,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{[i,j]}}}\Biggr{)},\quad\text{and}$
$D(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{2,i+1}+\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-1,j]}}}\Biggr{)},\quad\text{where}\quad Y_{\ell}=X_{2,\ell+1}-X_{1,\ell}.$
This map is a shifted version of the map studied in [CGS24]; see Appendix A for the precise connection. Note that we reverse the order of the inputs $\mathbf{W}_{1},\mathbf{W}_{2}$ in the application of $T$ and $D$. Also, note that the maps $\mathcal{P},T,D$ depend on the length $N$ of the periodicity; we will subsume dependence on this value in the notation and fix $N$, unless otherwise noted.
1.2. Polymers on $\mathbb{Z}^{2}$ and periodic environments
Definition 1.1.
Consider the 2-dimensional integer lattice $\mathbb{Z}^{2}$. We say that an up-right path of length $L\in\mathbb{N}$ is an ordered collection of points
$\pi:=\left\{(a_{i},b_{i})_{i\in\llbracket 1,L\rrbracket}\mid(a_{i+1},b_{i+1})-(a_{i},b_{i})\in\{(1,0),(0,1)\}\ \forall 1\leq i<L\right\}$ (1.3)
where $(a_{i},b_{i})\in\mathbb{Z}^{2}\ \forall i$ and the notation $\llbracket a,b\rrbracket$ denotes the set of $k\in\mathbb{Z}$ satisfying $a\leq k\leq b$. We also define the set of all up-right paths to be $\Pi_{\mathbb{Z}^{2}}$.
For a fixed path $\pi$, we say that $(a_{1},b_{1})$ is the initial point of the path while $(a_{L},b_{L})$ is the end point of the path. For given start and end points $(i,j),(n,m)\in\mathbb{Z}^{2}$, we define
$\Pi_{(i,j),(n,m)}:=\{\pi\in\Pi_{\mathbb{Z}^{2}}\mid(a_{1},b_{1})=(i,j)\ ,(a_{L},b_{L})=(n,m)\}$ (1.4)
to be the set of paths starting at $(i,j)$ and ending at $(n,m)$, noting $\Pi_{(i,j),(n,m)}\neq\varnothing\iff m\geq j$ and $n\geq i$. In this case, L=m+njiL=m+n-j-i for any $\pi$ in $\Pi_{(i,j),(n,m)}$. We say that two paths $\pi_{1},\pi_{2}\in\Pi_{\mathbb{Z}^{2}}$ are non-intersecting if $\pi_{1}\cap\pi_{2}=\varnothing$.
Also of interest to us are sets of non-intersecting paths on our graph. We define a multi-path $(\pi_{1},\ldots,\pi_{k})$ to be a non-intersecting $k$-tuple of paths. See Figure 2 for a visual example.
As in the case of single paths, we can consider the set of paths between given start and endpoints. For $k\in\mathbb{N}$ and $k$-tuples of ordered pairs $U=\bigl{(}(a_{i},b_{i})\bigr{)}_{i\in\llbracket 1,k\rrbracket}$, $V=\bigl{(}(c_{i},d_{i})\bigr{)}_{i\in\llbracket 1,k\rrbracket}$, we define the set of multi-paths between them to be
$\Pi_{U,V}:=\{\pi=(\pi_{1},\dots,\pi_{k})\mid\pi_{i}\in\Pi_{(a_{i},b_{i}),(c_{i},d_{i})}\ \forall i\in\llbracket 1,k\rrbracket,\ \pi_{i}\cap\pi_{j}=\varnothing\ \forall i\neq j\in\llbracket 1,k\rrbracket\}$ (1.5)
We will often require that there is a unique pairing of initial points in $U$ and end points in $V$ such that a non-intersecting multi-path exists. In other words, we require that if $U$ and $V$ each have $k$ elements, then there does not exist a nontrivial permutation $\sigma$ in the $k$th-order symmetric group $S_{k}$ such that $\Pi_{U,V_{\sigma}}\neq\varnothing$ where $V_{\sigma}:=\bigl{(}(c_{\sigma(i)},d_{\sigma(i)})\bigr{)}_{i\in\llbracket 1,k\rrbracket}$. We denote the pairs of $k$-tuples of ordered pairs satisfying this property as
$\Psi:=\bigl{\{}(U,V)\mid U,V\subseteq\mathbb{Z}^{2},\ |U|=|V|,\ \Pi_{U,V}\neq\varnothing,\ \text{ and }\Pi_{U,V_{\sigma}}=\varnothing\ \forall\sigma\in S_{|U|}\setminus\{\text{Id}\}\bigr{\}}$ (1.6)
[…]
Figure 2. Example multi-path in $\mathbb{Z}^{2}$ from $\{(0,0),(2,0)\}$ to $\{(2,4),(4,4)\}$
Definition 1.2.
For real-valued weights $\mathbf{W}=\bigl{(}\mathbf{W}_{i}\in\mathbb{R}^{\mathbb{Z}}\bigr{)}_{i\in\mathbb{Z}}=\bigl{(}W_{(i,j)}\in\mathbb{R}\mid i,j\in\mathbb{Z}\bigr{)}$ corresponding to the vertices of the directed lattice $\mathbb{Z}^{2}$, we define the polymer partition function with inverse temperature $\beta>0$ between the initial point $(i,j)$ and endpoint $(n,m)$ in $\mathbb{Z}^{2}$ as
$Z^{\beta,\mathbf{W}}(n,m\mid i,j):=\sum_{\pi\in\Pi_{(i,j),(n,m)}}\prod_{(r,\ell)\in\pi}e^{\beta W_{(r,\ell)}}.$ (1.7)
The partition function has a natural multi-path extension. For $(U,V)\in\Psi$, we define the multi-path partition function between $U$ and $V$ to be
$Z^{\beta,\mathbf{W}}(V\mid U):=\sum_{\pi\in\Pi_{U,V}}\prod_{(r,\ell)\in\pi}e^{\beta W_{(r,\ell)}}.$ (1.8)
We typically take $\beta=1$, in which case we omit the superscript $\beta$ in the notation.
Remark 1.3.
In Definition 1.2, we defined a polymer for arbitrary weights on $\mathbb{Z}^{2}$. In all results below (with the exception of Theorem 1.8) we specialize to the case where the weights are periodic in the vertical direction.
1.3. Invariance of periodic polymer partition functions
Let $(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ be the set of bi-infinite sequences of elements of $\mathbb{R}^{\mathbb{Z}_{N}}$. For $k\in\mathbb{Z}$, we define the operator $\mathcal{P}_{k}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ by its action:
$\mathcal{P}_{k}(\ldots,\mathbf{W}_{k-1},\mathbf{W}_{k},\mathbf{W}_{k+1},\mathbf{W}_{k+2},\ldots)=(\ldots,\mathbf{W}_{k-1},T(\mathbf{W}_{k+1},\mathbf{W}_{k}),D(\mathbf{W}_{k+1},\mathbf{W}_{k}),\mathbf{W}_{k+2},\ldots)$ (1.9)
Our first main result gives the invariance of the polymer partition functions under the iterated Pitman transform defined in (1.2).
Theorem 1.4.
Let $N\in\mathbb{N}$, and let $\mathbf{W}=(\mathbf{W}_{k}\in\mathbb{R}^{\mathbb{Z}_{N}}\mid k\in\mathbb{Z})=(W_{k,\ell}\mid k\in\mathbb{Z},\ell\in\mathbb{Z}_{N})$ be a collection of real-valued weights. Extend these periodically to weights on $\mathbb{Z}\times\mathbb{Z}$ by the condition $W_{k,j}=W_{k,\ell}$ whenever $j\equiv\ell\mod N$. If $k_{1},k_{2},\ldots,k_{m}\in\mathbb{Z}$ and the sets of initial points $U=\bigl{(}\bigl{(}a_{i},b_{i})_{i\in\llbracket 1,m\rrbracket}\bigr{)}$ and end points $V=\bigl{(}(c_{i},d_{i})_{i\in\llbracket 1,m\rrbracket}\bigr{)}$ satisfy $(U,V)\in\Psi$ and, for all $i,j\in\llbracket 1,m\rrbracket$, $a_{i}\neq k_{j}+1$ and $c_{i}\neq k_{j}$, then,
$Z^{\mathcal{P}_{k_{m}}\ldots\mathcal{P}_{k_{1}}\mathbf{W}}(V\mid U)=Z^{\mathbf{W}}(V\mid U).$ (1.10)
Remark 1.5.
See Figure 3 for a visual depiction of the invariance in Theorem 1.4. In Theorem 1.6 below, we show braid relations for the operators $\mathcal{P}_{k}$. Along with the fact that each $\mathcal{P}_{k}$ is an involution (see Lemma A.2), these relations imply that, if we have the equality of products of transpositions $(k_{m},k_{m}+1)(k_{m-1},k_{m-1}+1)\cdots(k_{1},k_{1}+1)=(k_{m}^{\prime},k_{m}^{\prime}+1)(k_{m-1}^{\prime},k_{m-1}^{\prime}+1)\cdots(k_{1}^{\prime},k_{1}^{\prime}+1)$, then $\mathcal{P}_{k_{m}}\ldots\mathcal{P}_{k_{1}}\mathbf{W}=\mathcal{P}_{k_{m}}^{\prime}\ldots\mathcal{P}_{k_{1}}^{\prime}\mathbf{W}$. Hence, we consider the action of $\mathcal{P}_{k_{m}}\ldots\mathcal{P}_{k_{1}}$ on $\mathbf{W}$ as the action of a general permutation, independent of its choice of decomposition. In this case, (1.11) gives an alternate condition for the set of endpoints whose partition function is preserved. An example is given in Figure 4 below. This connection is made precise in Lemma 4.2.
Figure 3. A depiction of the polymer invariance in Theorem 1.4. Each column is associated to a periodic vector of weights, $\mathbf{W}_{k}$. The horizontal lines are used to indicate a periodic environment for the polymer with $N=4$. In this example, we take $m=1$ and consider the action of the operator $\mathcal{P}_{3}$, which replaces the weights in columns $3$ and $4$ with $T(\mathbf{W}_{4},\mathbf{W}_{3})$ and $D(\mathbf{W}_{4},\mathbf{W}_{3})$, respectively. The partition function for each pair of points and the multi-path partition function for this collection of 33 pairs are both preserved under this transformation of the weights.
Theorem 1.4 is proved in Section 2. It is the first of its kind for a polymer in a periodic environment. The invariance under Pitman transformations in other contexts has been widely studied. The concept can be traced back to queuing theory from the works [Web79, TW87], where it was shown that the law of the departures process from two $\cdot/M/1$ queues in series is invariant under swapping of the service rates of the two queues. In the polymer context, this type of result can be traced back to the work of Noumi and Yamada [NY04], who gave a description of the geometric Robinson-Schendsted-Knuth correspondence in terms of certain matrix product relations. Corwin [Cor21] used this to give an analogous result to our Theorem 1.4 for a full-line version of the discrete Pitman transform. The proof of Theorem 1.4 likewise uses the machinery from [NY04], but we instead work with $\mathbb{Z}\times\mathbb{Z}$ periodic matrices instead of finite-dimensional matrices.
Theorem 1.4 limits to a zero-temperature analogue in Theorem 1.10. This zero-temperature limit is the periodic analogue of invariances shown in [DOV22, DNV22, DV24, BEM+25a]. The result in [DOV22] is particularly significant, as it was key in constructing the directed landscape, the central limiting object of the KPZ universality class.
It would be interesting to see if Theorem 1.4 could be the starting point for constructing and proving convergence to the periodic directed landscape, but we note that the action of this transform is qualitatively different than that in [DOV22]. Specifically, in the full-line case, the action of the iterated Pitman transforms and independent collection of Brownian motions and transforms them to the Brownian melon–a set of nonintersecting paths whose geodesic structure is more amenable to asymptotic analysis. By contrast, in this periodic setting when we take the weights to be i.i.d. with the inverse-gamma distribution, the transformed weights are equal in distribution to the original weights. This is the Burke property shown in [CGS24, Proposition 5.12], and extended to an inhomogeneous setting in Proposition 4.1. As a corollary, we obtain a permutation invariance result for the multi-path inhomogeneous inverse-gamma polymer below in Theorem 1.7. This extends the previous result of [BEM+25a, Theorem 3.1] to a positive-temperature and multi-path setting. In this sense, the periodic Pitman transform is more analogous to a limiting case of the full-line Pitman transform, where the initial point of a random walk is sent to -$\infty$ and the output of the transform is recentered at 0 (see [DNV22, DV24]). This limiting full-line setting is where we see analogous Burke properties, as in [OY01, OY02].
1.4. Braid relations
We now describe our next result, which gives a braid relation for the Pitman operators defined in (1.9). The next theorem gives a braid relation for these operators, which is the same as that satisfied for the full-line Pitman transform [BBO05]. See also [DV24]. Combined with the fact that the map $\mathcal{P}_{k}$ is an involution (see Lemma A.2), the operators $\mathcal{P}_{k}$ generate a representation of the infinite symmetric group (note that the same braid relations are satisfied by the nearest-neighbor transpositions $(k,k+1)$).
Theorem 1.6.
For $k\in\mathbb{Z}$,
$\mathcal{P}_{k}\mathcal{P}_{k+1}\mathcal{P}_{k}=\mathcal{P}_{k+1}\mathcal{P}_{k}\mathcal{P}_{k+1}.$
Theorem 1.6 is proved in Section 3. The key is the matrix relation in Corollary 2.5 below, which is also used to prove Theorem 1.4. When written out, the braid relation consists of three equalities ((3.1)-(3.3) below). The third of these equalities was previously proved in [CGS24, Proposition 5.13]. With that as our starting point, we iterate the matrix relation in Corollary 2.5 in two different ways to obtain and use the interpretation of the matrix product as the polymer partition function to obtain the other two equalities. Specifically, Lemma 3.1 is the key innovation that allows us to complete the proof.
1.5. Permutation invariance
In this section, we consider an inhomogeneous, periodic inverse-gamma polymer. Let $\lambda>0$ and $\mathbf{a}\in\mathbb{R}^{\mathbb{Z}}$ and $\mathbf{b}\in\mathbb{R}^{\mathbb{Z}_{N}}$ be sequences of real numbers such that $a_{i}+b_{j}>0$ for $i\in\mathbb{Z}$ and $j\in\mathbb{Z}_{N}$. We take independent weights $\mathbf{W}^{\mathbf{a},\mathbf{b}}=\bigl{(}W_{(i,j)}\mid i\in\mathbb{Z},\ j\in\mathbb{Z}_{N}\bigr{)}$ such that $W_{(i,j)}\sim\text{Log-Inv-Gamma}(a_{i}+b_{j},\lambda)$ for $i\in\mathbb{Z}$ and $j\in\mathbb{Z}_{N}$ (see Definition 1.15 below). Extend these periodically to weights on $\mathbb{Z}\times\mathbb{Z}$. We call a permutation $\sigma:\mathbb{Z}\to\mathbb{Z}$ a finite permutation if $\sigma$ fixes all but finitely many coordinates. For a finite permutation $\sigma$, we let $\mathcal{U}^{\sigma}$ be the set of tuples $(U,V)\in\Psi$, such that
[…]
Theorem 1.7.
For vectors $\mathbf{a}\in\mathbb{R}^{\mathbb{Z}}$ and $\mathbf{b}\in\mathbb{R}^{\mathbb{Z}_{N}}$ satisfying $a_{i}+b_{j}>0$ for $i\in\mathbb{Z}$ and $j\in\mathbb{Z}_{N}$, define periodic weights $\mathbf{W}^{\mathbf{a},\mathbf{b}}$ as above and extend them to $\mathbb{Z}\times\mathbb{Z}$. Then, for any finite permutation $\sigma:\mathbb{Z}\to\mathbb{Z}$, we have
$\Bigl{(}Z^{\mathbf{W}^{\mathbf{a},\mathbf{b}}}(V\mid U):(U,V)\in\mathcal{U}^{\sigma}\Bigr{)}\overset{d}{=}\Bigl{(}Z^{\mathbf{W}^{\sigma\mathbf{a},\mathbf{b}}}(V\mid U):(U,V)\in\mathcal{U}^{\sigma}\Bigr{)},$
where $\sigma\mathbf{a}=(a_{\sigma(i)})_{i\in\mathbb{Z}}$.
The condition (1.11) ensures that, between the pairs of points in $(U,V)$, the column rates are either permuted among themselves or are left unchanged. See Figure 4 for an example.
Figure 4. An example of an element $(U,V)\in\mathcal{U}^{\sigma}$, where $U$ and $V$ each contain 33 elements. The gray horizontal lines are used to indicate a periodic environment with $N=4$. In this figure, the column parameters of columns $3,5$, and 66 (blue) are permuted, and the column parameters in columns $8$ and 99 (red) are permuted. The condition is that, for each of the three endpoints $(x,y)$ in $U$ (lower left), $\sigma(\mathbb{Z}_{<x})=\mathbb{Z}_{<x}$, while for each of the three endpoints $(x,y)$ in $V$ (upper right), $\sigma(\mathbb{Z}_{>x})=\mathbb{Z}_{>x}$. The distributional equality in Theorem 1.7 holds jointly over all endpoint pairs $(U,V)\in\Psi$ satisfying this property.
We can allow for permuting the $\mathbf{b}$ parameters as well, with doubly periodic weights $(W_{(i,j)}\mid i\in\mathbb{Z}_{M},j\in\mathbb{Z}_{N})$ for $M,N\in\mathbb{N}$ when we restrict to points that lie inside the box $\llbracket 0,M\rrbracket\times\llbracket 0,N\rrbracket$. In this case, the model isn’t actually periodic at all; it is just a restriction of the full-space polymer to a finite grid. For finite permutations $\sigma,\tau:\mathbb{Z}\to\mathbb{Z}$, define the set $\mathcal{U}^{\sigma,\tau}$ to be the set of tuples $(U,V)$, where $U,V$ are each sets of endpoint pairs in $\mathbb{Z}\times\mathbb{Z}$ satisfying (1.11) as well as
[…]
The following is an extension of the recent result [BEM+25a, Theorem 3.1] to multi-paths and to the positive-temperature case. A sketch of the proof for the extension to multi-paths in zero temperature was previously discussed in [BEM+25a, Remark 4.3]. The technique follows the same outline as developed in [BEM+25a], as we use a coupling of the weights given by applying a Pitman transform to neighboring rows and columns. However, the coupling we use in this proof is different because we treat the weights as periodic outside a large box and apply the periodic Pitman transform. The algebra involved in this positive-temperature case is also different.
Theorem 1.8.
For vectors $\mathbf{a}\in\mathbb{R}^{\mathbb{Z}}$ and $\mathbf{b}\in\mathbb{R}^{\mathbb{Z}}$ satisfying $a_{i}+b_{j}>0$ for $i,j\in\mathbb{Z}$ define independent weights $\mathbf{W}^{\mathbf{a},\mathbf{b}}=\bigl{(}W_{(i,j)}:i,j\in\mathbb{Z}\bigr{)}$ such that $W_{(i,j)}\sim\text{Log-Inv-Gamma}(a_{i}+b_{j},\lambda)$ for all $i,j$. Then, for any finite permutations $\sigma,\tau:\mathbb{Z}\to\mathbb{Z}$, we have
$\Bigl{(}Z^{\mathbf{W}^{\mathbf{a},\mathbf{b}}}(V\mid U):(U,V)\in\mathcal{U}^{\sigma,\tau}\Bigr{)}\overset{d}{=}\Bigl{(}Z^{\mathbf{W}^{\sigma\mathbf{a},\tau\mathbf{b}}}(V\mid U):(U,V)\in\mathcal{U}^{\sigma,\tau}\Bigr{)},$
where $\sigma\mathbf{a}=(a_{\sigma(i)})_{i\in\mathbb{Z}}$ and $\tau\mathbf{b}=(b_{\tau(j)})_{j\in\mathbb{Z}}$.
Theorems 1.7 and 1.8 are proved in Section 4. When the initial point is fixed, invariances of this flavor can be deduced from the symmetry of Schur functions and Whittaker functions (see, for example, [BP08, COSZ14, OSZ14, BC14]). A similar type of invariance for last-passage percolation was discovered in the work of Dauvergne [Dau22, Theorem 1.5], motivated by a shift-invariance result previously proved in Borodin, Gorin, and Wheeler [BGW22]. Connections between these results are discussed further in [BEM+25a, Section 3].
1.6. Zero-temperature limits
We also define the zero-temperature discrete periodic Pitman transform as the map $\widetilde{\mathcal{P}}:\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}\to\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$ defined by
$\widetilde{\mathcal{P}}(\mathbf{W}_{1},\mathbf{W}_{2})$ $=\bigl{(}\widetilde{T}(\mathbf{W}_{2},\mathbf{W}_{1}),\widetilde{D}(\mathbf{W}_{2},\mathbf{W}_{1})\bigr{)},\quad\text{where}$ (1.13)
$\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{1,i-1}+\max_{j\in\mathbb{Z}_{N}}Y_{[i-1,j]}-\max_{j\in\mathbb{Z}_{N}}Y_{[i,j]},\quad\text{and}$
$\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{2,i+1}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]},\quad\text{where}\quad Y_{\ell}=X_{2,\ell+1}-X_{1,\ell}.$
The mapping $\widetilde{D}$ is similar to, but not equivalent to a mapping used to describe the invariant measures for multi-type TASEP by Ferrari and Martin [FM07]. See [CGS24, Appendix F] for more on this connection. One immediately sees that we can obtain $\widetilde{\mathcal{P}}$ as a limiting case of the positive-temperature discrete periodic Pitman transform:
$\widetilde{\mathcal{P}}(\mathbf{W}_{1},\mathbf{W}_{2})=\lim_{\beta\to\infty}\frac{1}{\beta}\mathcal{P}(\beta\mathbf{W}_{1},\beta\mathbf{W}_{2})$
Furthermore, this convergence is uniform as a function of the vectors $\mathbf{W}_{1},\mathbf{W}_{2}$. See Lemma A.1 for details.
Definition 1.9.
For real-valued weights $\mathbf{W}=\bigl{(}W_{(i,j)}\in\mathbb{R}\mid i,j\in\mathbb{Z}\bigr{)}$ corresponding to the vertices of the directed lattice $\mathbb{Z}^{2}$, we define the last-passage percolation (LPP) time for both a single pair of points and multi-paths analogously as in (1.7)-(1.8):
$G^{\mathbf{W}}(V\mid U)=\max_{\pi\in\Pi_{U,V}}\sum_{(r,\ell)\in\pi}W_{(r,\ell)}.$
For $k\in\mathbb{Z}$, define the map $\widetilde{\mathcal{P}}_{k}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ by (1.9), with $T,D$ replaced by $\widetilde{T},\widetilde{D}$.
With this notation defined, we get the following zero-temperature analogues of our main theorems.
Theorem 1.10.
The same invariance in Theorem 1.4 holds when $Z^{\mathbf{W}}$ is replaced by $G^{\mathbf{W}}$, and the maps $\mathcal{P}_{k_{j}}$ are replaced by $\widetilde{\mathcal{P}}_{k_{j}}$.
Theorem 1.11.
The same braid relations in Theorem 1.6 hold for the maps $\widetilde{\mathcal{P}}_{k}$.
Theorem 1.12.
The analogues of Theorems 1.7 and 1.8 hold, when $Z^{\mathbf{W}}$ is replaced by $G^{\mathbf{W}}$ and, instead of log-inverse-gamma weights, the weights satisfy either $W_{(i,j)}\sim\text{Geom}(q_{i}r_{j})$ for all $i,j$, where $q_{i}r_{j}\in(0,1)$, or $W_{(i,j)}\sim\text{Exp}(a_{i}+b_{j})$ for all $i,j$, where $a_{i}+b_{j}>0$ (see Definitions 1.13-1.14 below).
The details of the proofs of these theorems are given in Appendix B.
1.7. Organization of the paper
Section 2 is focused on proving the partition function invariance in Theorem 1.4. The braid relations in Theorem 1.6 are proved in Section 3. Section 4 proves an inhomogeneous Burke property for the periodic Pitman transform and uses it to prove Theorems 1.7 and 1.8. Appendix A contains some prior results from the work [CGS24] and some auxiliary technical results. The necessary details for the zero-temperature analogues in Theorems 1.10-1.12 are proved in Appendix B.
1.8. Notation and conventions
We use $\overset{d}{=}$ to indicate that two random variables are equal in distribution. Below, we define our notation for geometric, exponential, and log-inverse-gamma random variables.
Definition 1.13 (Geometric Random Variable).
A random variable $X\in\mathbb{Z}_{\geq 1}$ has a geometric distribution with parameter $q\in[0,1]$ if for all $n\in\mathbb{Z}_{\geq 0}$, we have that
$\mathbb{P}(X=n)=q^{n}(1-q)$
In this case, we write $X\sim\text{Geom}(q)$.
Definition 1.14 (Exponential Random Variable).
A random variable $X\in\mathbb{R}$ has an exponential distribution with rate $\alpha>0$ if
$\mathbb{P}(X\in dx)=\begin{cases}\alpha e^{-\alpha x}&x\geq 0\\ 0&x<0\end{cases}$
In this case, we write $X\sim\text{Exp}(\alpha)$.
Definition 1.15 (Log-Inverse-Gamma Random Variable).
A random variable $X\in\mathbb{R}$ has a log-inverse-gamma distribution with shape $\gamma>0$ and scale $\lambda>0$ if
$\mathbb{P}(X\in dx)=\frac{\lambda^{\gamma}}{\Gamma(\gamma)}e^{-\gamma x}e^{-\lambda e^{-x}}$
In this case, we write $X\sim\text{Log-Inv-Gamma}(\gamma,\lambda)$.
1.9. Acknowledgements
This work was completed during the Columbia Math Summer Undergraduate Research Program in Summer 2025, and we acknowledge support for this program from the Columbia University math department. We thank George Dragomir for organizing this program. During the program, E.S. and X.Z. were partially supported by Ivan Corwin’s Simons Investigator Grant #929852. In addition, X.Z. was supported by Ivan Corwin’s NSF grant DMS:2246576. E.R.E. was supported by Princeton University’s Office of Undergraduate Research OURSIP Program through the Frances Lane Summer Research Fund. We also thank Ivan Corwin and Amol Aggarwal for helpful discussions related to this project.
2. Invariance of polymer partition function
In this section, we prove Theorem 1.4. The key to our proof is a matrix encoding of polymer partition functions, analogous to that developed in [NY04]. Section 2.1 describes the necessary algebraic properties of these matrices, and then Section 2.2 makes the connection to the polymer partition function and completes the proof.
2.1. Matrix encoding
In order to prove our polymer partition function invariance, we follow the matrix encoding method introduced at Noumi and Yamada in [NY04]. The key difference is that we will work with infinite matrices.
Definition 2.1.
For $\mathbf{x}\in\mathbb{R}^{\mathbb{Z}}$, we introduce two functions $E:\mathbb{R}^{\mathbb{Z}}\to\mathbb{R}^{\mathbb{Z}\times\mathbb{Z}}$ and $H:\mathbb{R}^{\mathbb{Z}}\to\mathbb{R}^{\mathbb{Z}\times\mathbb{Z}}$ as follows:
$E(\mathbf{x}):=\sum_{i\in\mathbb{Z}}e^{x_{i}}E_{i,i}+\sum_{i\in\mathbb{Z}}E_{i,i+1}\,,\quad H(\mathbf{x}):=\sum_{i\leq j}e^{x_{i}+x_{i+1}+\ldots+x_{j}}E_{i,j}$ (2.1)
where $E_{p,q}$ is the $\mathbb{Z}\times\mathbb{Z}$ matrix with a $1$ at index $(p,q)$ and 0s elsewhere.
These matrix functions admit natural extensions for $\mathbf{x}\in\mathbb{R}^{\mathbb{Z}_{N}}$, where, for $k\notin\{1,\ldots,N\}$, we simply set $x_{k}\coloneqq x_{\ell}$, where $\ell$ is the unique element of $\{1,\ldots,N\}$ satisfying $\ell\equiv k\pmod{N}$. All of the same properties that hold for $\mathbb{Z}$ trivially also hold for $\mathbb{Z}_{N}$, via the inclusion map $i:\mathbb{R}^{\mathbb{Z}_{N}}\to\mathbb{R}^{\mathbb{Z}}$.
The intuition behind this approach is that matrix multiplication $H(\mathbf{x}_{1})\cdots H(\mathbf{x}_{n})$ naturally encodes the partition function across a certain diagram with vertex weights given by the arrays $\mathbf{x}_{1},\ldots,\mathbf{x}_{n}$ as discussed in Lemma 2.6. Meanwhile, the outputs of map $E$ are more algebraically straightforward and closely related to the matrix inverse of $H$, which we prove in the next Lemma 2.2 as an infinite-dimensional analogue of [NY04, Equation 1.5].
Lemma 2.2.
For all $\mathbf{x}\in\mathbb{R}^{\mathbb{Z}}$, we have
$\Delta H(\mathbf{x})\Delta E(-\mathbf{x})=E(-\mathbf{x})\Delta H(\mathbf{x})\Delta=\text{Id}_{\mathbb{Z}},$ (2.2)
where $\Delta=\operatorname{diag}((-1)^{i-1})_{i\in\mathbb{Z}}$ and $\text{Id}_{\mathbb{Z}}=\operatorname{diag}(1)_{i\in\mathbb{Z}}$.
Proof.
We compute $H(\mathbf{x})\Delta E(-\mathbf{x})$ as follows:
$H(\mathbf{x})\Delta E(-\mathbf{x})$ $=H(\mathbf{x})\left(\sum_{j\in\mathbb{Z}}(-1)^{j-1}e^{-x_{j}}E_{j,j}+\sum_{j\in\mathbb{Z}}(-1)^{j-1}E_{j,j+1}\right)$
$=H(\mathbf{x})\sum_{j\in\mathbb{Z}}(-1)^{j-1}e^{-x_{j}}E_{j,j}+H(\mathbf{x})\sum_{j\in\mathbb{Z}}(-1)^{j-1}E_{j,j+1}$
$=\sum_{i\in\mathbb{Z}}(-1)^{i-1}E_{i,i}+\sum_{i\leq j}\bigl{(}(-1)^{j-1}e^{x_{i}+\cdots+x_{j-1}}+(-1)^{j-2}e^{x_{i}+\cdots+x_{j-1}}\bigr{)}E_{i,k}$
$=\Delta\ ,$
where the third equality is obtained by shifting the index of the second summand from $j$ to j1j-1 and last equality above follows by a telescoping of terms. A similar calculation shows that $E(-\mathbf{x})\Delta H(\mathbf{x})=\Delta$. This concludes the proof, since one can easily check that $\Delta^{2}=Id$. ∎
Remark 2.3.
Noticing that $\Delta H(\mathbf{x})\Delta$ is a left and right inverse for $E(-\mathbf{x})$, we see that it is an inverse in the semi-group of finite valued upper-triangular $\mathbb{Z}\times\mathbb{Z}$ matrices under matrix multiplication. Under this notation, we see that Lemma 2.2 translates to
$H(\mathbf{x})=\Delta E(-\mathbf{x})^{-1}\Delta.$ (2.3)
Lemma 2.4.
For $\mathbf{W}_{1},\mathbf{W}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}$, we have that
$E(-\mathbf{W}_{2})E(-\mathbf{W}_{1})=E(-D(\mathbf{W}_{2},\mathbf{W}_{1}))E(-T(\mathbf{W}_{2},\mathbf{W}_{1})).$ (2.4)
Proof.
By direct computation and applying Lemma A.5 in the third equality below, we see that
$\quad\,E(-D(\mathbf{W}_{2},\mathbf{W}_{1}))E(-T(\mathbf{W}_{2},\mathbf{W}_{1}))$
$=\left(\sum_{i\in\mathbb{Z}}e^{-D_{i}(\mathbf{W}_{2},\mathbf{W}_{1})}E_{i,i}+\sum_{i\in\mathbb{Z}}E_{i,i+1}\right)\left(\sum_{i\in\mathbb{Z}}e^{-T_{i}(\mathbf{W}_{2},\mathbf{W}_{1})}E_{i,i}+\sum_{i\in\mathbb{Z}}E_{i,i+1}\right)$
$=\sum_{i\in\mathbb{Z}}e^{-(D_{i}(\mathbf{W}_{2},\mathbf{W}_{1})+T_{i}(\mathbf{W}_{2},\mathbf{W}_{1}))}E_{i,i}+\sum_{i\in\mathbb{Z}}\left(e^{-D_{i}(\mathbf{W}_{2},\mathbf{W}_{1})}+e^{-T_{i+1}(\mathbf{W}_{2},\mathbf{W}_{1})}\right)E_{i,i+1}+\sum_{i\in\mathbb{Z}}E_{i,i+2}$
$=\sum_{i\in\mathbb{Z}}e^{-(W_{1,i}+W_{2,i})}E_{i,i}+\sum_{i\in\mathbb{Z}}\left(e^{-W_{1,i+1}}+e^{-W_{2,i}}\right)E_{i,i+1}+\sum_{i\in\mathbb{Z}}E_{i,i+2}$
$=\left(\sum_{i\in\mathbb{Z}}e^{-W_{2,i}}E_{i,i}+\sum_{i\in\mathbb{Z}}E_{i,i+1}\right)\left(\sum_{i\in\mathbb{Z}}e^{-W_{1,i}}E_{i,i}+\sum_{i\in\mathbb{Z}}E_{i,i+1}\right)$
$=E(-\mathbf{W}_{2})E(-\mathbf{W}_{1}).\qed$
Corollary 2.5.
For $\mathbf{W}_{1},\mathbf{W}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}$, we have
$H(\mathbf{W}_{1})H(\mathbf{W}_{2})=H(T(\mathbf{W}_{2},\mathbf{W}_{1}))H(D(\mathbf{W}_{2},\mathbf{W}_{1})).$ (2.5)
Proof.
First, we justify that this matrix product has finite entries. Because these $H$ matrices (as well as the $E$ matrices) are upper triangular, each individual index in the product consists of a finite sum of products of elements in the respective $H$ matrices. While each row and column has infinitely many non-zero terms, there are only a finite number of non-zero terms in the dot product. Further, the product is upper triangular, so by repeating this argument, we can see that an arbitrary product of $H$ (or $E$) matrices is finite valued.
With this in mind, the proof follows from Lemma 2.2, since:
[…]
$=\Delta E(-\mathbf{W}_{1})^{-1}E(-\mathbf{W}_{2})^{-1}\Delta$
$=\Delta\bigl{(}E(-\mathbf{W}_{2})E(-\mathbf{W}_{1})\bigr{)}^{-1}\Delta$
[…]
$=(\Delta E(-T(\mathbf{W}_{2},\mathbf{W}_{1}))^{-1}\Delta)(\Delta E(-D(\mathbf{W}_{2},\mathbf{W}_{1}))^{-1}\Delta)$
$=H(T(\mathbf{W}_{2},\mathbf{W}_{1}))H(D(\mathbf{W}_{2},\mathbf{W}_{1})).\qed$
2.2. Relation to the periodic polymer
We now show the following relation between the $H$ matrices and the polymer partition function:
Lemma 2.6.
Let $\mathbf{W}=(\mathbf{W}_{i})_{i\in\mathbb{Z}}\in(\mathbb{Z}_{N})^{\mathbb{Z}}$ and $k,\ell\in\mathbb{Z}$ such that $\ell\geq k$, then:
$\Bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})\Bigr{)}_{(i,j)}=Z^{\mathbf{W}}(\ell,j\mid k,i)\quad\forall i,j\in\mathbb{Z}.$
In other words, the partition function across any two points $(k,i)$ and $(\ell,j)$ on the directed lattice $\mathbb{Z}^{2}$ with periodic weights $\mathbf{W}$ is given by the $(i,j)$th entry in the matrix product of the $H$ matrices.
Proof.
We prove the result by induction on $\ell\geq k$. For the base case, $\ell=k$, by definition we have that
$H(\mathbf{W}_{k})_{(i,j)}=\begin{cases}\exp\Biggl{(}{\sum\limits_{r=i}^{j}}W_{(k,r)}\Biggr{)},\quad&j\geq i\\ 0,\ &\text{otherwise}\end{cases}$
Similarly, on a single column, we either have a unique up/right path between $(k,i)$ and $(k,j)$ picking up all the weights between $i$ and $j$ or have no path between them, depending on whether $j\geq i$. Thus,
$Z^{W}(k,j\mid k,i)=\begin{cases}\exp\Biggl{(}{\sum\limits_{r=i}^{j}}W_{(k,r)}\Biggr{)},\quad&j\geq i\\ 0,\ &\text{otherwise}\end{cases}=H(\mathbf{W}_{k})_{(i,j)}$
This proves our base case. Now, let us assume that it holds for some $\ell$, and prove it holds for $\ell+1$. We split into two cases: $j\geq i$ and $j<i$.
For $j<i$, notice that $H$ is upper triangular, so their product is as well. Thus, the $(i,j)$-th index of our matrix product is 0 for $j<i$. This matches up with our partition function, as there are no up/right paths whose initial point is lower than its end point.
For the case $j\geq i$, we notice that every path from $(k,i)$ to $(\ell+1,j)$ consists of a path from $(k,i)$ to $(\ell,n)$ concatenated with a path from $(\ell+1,n)$ to $(\ell+1,j)$, for some $i\leq n\leq j$. Thus,
$Z^{\mathbf{W}}(\ell+1,j\mid k,i)$ $=\sum\limits_{r=i}^{j}\Biggl{(}Z^{\mathbf{W}}(\ell,r\mid k,i)\exp\bigl{(}\sum_{s=r}^{j}W_{(\ell+1,s)}\bigr{)}\Biggr{)}$
$=\sum\limits_{r=i}^{j}\Biggl{(}Z^{\mathbf{W}}(\ell,r\mid k,i)\Bigl{(}H(\mathbf{W}_{\ell+1})\Bigr{)}_{(r,j)}\Biggr{)}.$
By our inductive hypothesis, we know that this equals
$\sum\limits_{r=i}^{j}\Biggl{(}\Bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})\Bigr{)}_{(i,r)}\Bigl{(}H(\mathbf{W}_{\ell+1})\Bigr{)}_{(r,j)}\Biggr{)}.$
Remembering now that $\bigl{(}H(\mathbf{W}_{\ell+1})\bigr{)}_{(r,j)}=0$ for $r>j$, and $\bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})\bigr{)}_{(i,r)}=0$ for $r<i$, we see that
$\quad\;Z^{\mathbf{W}}(\ell+1,j\mid k,i)$
$=\sum\limits_{r=-\infty}^{\infty}\Biggl{(}\Bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})\Bigr{)}_{(i,r)}\Bigl{(}H(\mathbf{W}_{\ell+1})\Bigr{)}_{(r,j)}\Biggr{)}$
$=\Bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})H(\mathbf{W}_{\ell+1})\Bigr{)}_{(i,j)}.\qed$
We now show the invariance of polymer partition function under the periodic Pitman transform. Retaining the periodic weight vectors $\mathbf{W}=(\mathbf{W}_{r}\in\mathbb{R}^{\mathbb{Z}_{N}}:r\in\mathbb{Z})$ as before, for $k\in\mathbb{Z}$, recall the operator $\mathcal{P}_{k}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ where
$\mathcal{P}_{k}\mathbf{W}=(\ldots,\mathbf{W}_{k-1},T(\mathbf{W}_{k+1},\mathbf{W}_{k}),D(\mathbf{W}_{k+1},\mathbf{W}_{k}),\mathbf{W}_{k+2,\ldots}).$
Proposition 2.7.
Fix $a,c\in\mathbb{Z}$ such that $a\leq c$, and let $k_{1},\dots,k_{m}\in\mathbb{Z}$ such that for all $i\in\llbracket 1,m\rrbracket$, $a\neq k_{i}+1$ and $c\neq k_{i}$. Then, for all $b,d\in\mathbb{Z}$,
$Z^{\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}\mathbf{W}}(c,d\mid a,b)=Z^{\mathbf{W}}(c,d\mid a,b).$
Proof.
We will prove this by induction on $m$. By Lemma 2.6,
$Z^{\mathbf{W}}(c,d\mid a,b)=\bigl{(}H(\mathbf{W}_{a})H(\mathbf{W}_{a+1})\cdots H(\mathbf{W}_{c})\bigr{)}_{(b,d)},$
which we claim is preserved by the action $\mathcal{P}_{k_{m}}\ldots\mathcal{P}_{k_{1}}$ for any $b,d\in\mathbb{Z}$. In our base case $n=1$, we split into three sub-cases: $k_{1}<a-1$, $k_{1}>c$, and $a\leq k_{1}<c$. In the first two sub-cases, $\mathbf{W}_{a}\ldots\mathbf{W}_{c}$ are unaffected by $\mathcal{P}_{k_{1}}$, so our invariance holds trivially.
If $a\leq k_{1}<c$, then Corollary 2.5 gives that
$Z^{\mathbf{W}}(c,d\mid a,b)$ $=\bigl{(}H(\mathbf{W}_{a})H(\mathbf{W}_{a+1})\ldots H(\mathbf{W}_{k_{1}})H(\mathbf{W}_{k_{1}+1})\ldots H(\mathbf{W}_{c})\bigr{)}_{(b,d)}$
[…]
$=Z^{\mathcal{P}_{k_{1}}\mathbf{W}}(c,d\mid a,b).$
Now, we assume by way of induction that the invariance holds for m1m-1 transforms. Again, if $k_{m}<a-1$ or $k_{m}>c$ our equivalence holds trivially. Otherwise, let $\mathbf{W}^{\prime}=\mathcal{P}_{k_{m-1}}\ldots\mathcal{P}_{k_{1}}\mathbf{W}$. Then
$Z^{\mathbf{W}}(c,d\mid a,b)$ $=Z^{\mathbf{W}^{\prime}}(c,d\mid a,b)$
$=Z^{\mathcal{P}_{k_{m}}\mathbf{W}^{\prime}}(c,d\mid a,b)$
$=Z^{\mathcal{P}_{k_{m}}\mathcal{P}_{k_{m-1}}\ldots\mathcal{P}_{k_{1}}\mathbf{W}}(c,d\mid a,b).$
This concludes our proof, noting that the second equality follows from the case $m=1$. ∎
With Proposition 2.7 proved, we now prove Theorem 1.4 by extending to the multi-path result.
Proof of Theorem 1.4..
We use the Lindström-Gessel-Viennot (LGV) lemma (cf. [AZ18, Chapter 32], originally from [Lin73, GV85]) to prove the theorem. Because LGV applies only to edge-weighted graphs, we need to transfer our vertex weights to edge weights in the following way: for each vertex $(r,s)$, split it into $(r,s)_{\text{in}}$ and $(r,s)_{\text{out}}$, with edge $\bigl{(}(r,s)_{\text{in}},(r,s)_{\text{out}}\bigr{)}$ given the weight $\omega_{(r,s)}:=e^{W_{(r,s)}}$, and edges $\bigl{(}(r,s)_{\text{out}},(r,s+1)_{\text{in}}\bigr{)}$ and $\bigl{(}(r,s)_{\text{out}},(r+1,s)_{\text{in}}\bigr{)}$ given the weight 1. See Figure 5. Any path which started at $(r,s)$ starts at $(r,s)_{\text{in}}$, while any path ending at $(r,s)$ ends at $(r,s)_{\text{out}}$. Under this graph transformation, our partition function on the pre-transformed graph equals the sum over weight products along possible paths on the transformed diagram, so we can apply LGV.
[…]
Figure 5. Local transformation of a vertex-weighted node (left) with weight $W_{(r,s)}$ into an edge-weighted directed graph (right), where the internal edge carries weight $\omega_{(r,s)}=\exp(W_{(r,s)})$ and external edges carry weight 1. Paths in the original graph correspond to paths from “in” nodes to “out” nodes in the transformed graph, preserving path weights.
Write the path matrix $M=\bigl{(}Z^{\mathbf{W}}(c_{j},d_{j}\mid a_{i},b_{i})\bigr{)}_{(i,j)\in\llbracket 1,m\rrbracket}$. By LGV, we have
$\det M=\sum\limits_{P=(P_{1}\ldots P_{m})\in\Pi_{U,V_{\sigma}}\ \forall\sigma\in S_{m}}\operatorname{sgn}(\sigma)\prod\limits_{(i,j)\in P}e^{W_{(i,j)}}.$
By definition of $\Psi$, for $\sigma\neq\text{Id}$, $\Pi_{U,V_{\sigma}}=\varnothing$. Because $\operatorname{sgn}(\text{Id})=1$, we have that
$\det M=\sum\limits_{P=(P_{1}\ldots P_{m})\in\Pi_{U,V}}\prod\limits_{(i,j)\in P}e^{W_{(i,j)}}=Z^{\mathbf{W}}(V\mid U).$
By Proposition 2.7, each $M_{i,j}=Z^{\mathbf{W}}(c_{j},d_{j}\mid a_{i},b_{i})$ remains invariant under $\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}$, so $M$ is unchanged, and so $\det M$ is invariant. Thus, $Z^{\mathbf{W}}(V\mid U)=\det M$ is invariant. ∎
3. Braid relations: proof of Theorem 1.6
In this section, we prove the braid relation stated in Theorem 1.6. Notice that this relation translates into the following three equalities:
$T\Bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\biggr{)}=T\Bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\Bigr{)},$ (3.1)
$D\Bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\Bigr{)}=T\Bigl{(}D(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),D\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)}\Bigr{)},$ (3.2)
$D\Bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\Bigr{)}=D\Bigl{(}D(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),D\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)}\Bigr{)}.$ (3.3)
Equation (3.3) follows from [CGS24, Proposition 5.13], as explained in Lemma A.8 of the appendix. We now use the partition function invariance, Theorem 1.4, to demonstrate that (3.1) implies both (3.2) and (3.3). The main tool will be the following lemma:
Lemma 3.1.
For $\mathbf{Y}=(\mathbf{Y}_{i})_{i\in\mathbb{Z}},\mathbf{Y}^{\prime}=(\mathbf{Y}^{\prime}_{i})_{i\in\mathbb{Z}}\in(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ satisfying $\mathfrak{s}(\mathbf{Y}_{1})=\mathfrak{s}(\mathbf{Y}^{\prime}_{1})$, $\mathfrak{s}(\mathbf{Y}_{2})=\mathfrak{s}(\mathbf{Y}_{2}^{\prime})$, we have that $H(\mathbf{Y}_{1})H(\mathbf{Y}_{2})=H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}_{2}^{\prime})$ if and only if $\mathbf{Y}_{1}=\mathbf{Y}^{\prime}_{1},\mathbf{Y}_{2}=\mathbf{Y}_{2}^{\prime}$.
Before proving Lemma 3.1, we first use Lemmas A.8 and 3.1 to complete the proof of Theorem 1.6:
Proof of Theorem 1.6..
Using the partition function invariance under the operator $\mathcal{P}_{i+1}\mathcal{P}_{i}\mathcal{P}_{i+1}$ yields:
$H\bigl{(}\mathbf{W}_{k}\bigr{)}H\bigl{(}\mathbf{W}_{k+1}\bigr{)}H\bigl{(}\mathbf{W}_{k+2}\bigr{)}=$
[…]
$\qquad H\Bigl{(}D\bigl{(}D(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),D\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)}\bigr{)}\Bigr{)}.$ (3.4)
On the other hand, we can also apply the partition function invariance under the operator $\mathcal{P}_{i}\mathcal{P}_{i+1}\mathcal{P}_{i}$:
$H\bigl{(}\mathbf{W}_{k}\bigr{)}H\bigl{(}\mathbf{W}_{k+1}\bigr{)}H\bigl{(}\mathbf{W}_{k+2}\bigr{)}=$
[…]
$\qquad H\Bigl{(}D\bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)}\Bigr{)}H\Bigl{(}D(\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k}))\Bigr{)}.$ (3.5)
From Lemma A.8, we know that the last terms in (3.4) and (3.5) are equal. Because $H$ is invertible (see Lemma 2.2), we can equate both equations and cancel out the last term, obtaining that:
$H\Bigl{(}T\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)}\Bigr{)}H\Bigl{(}T\bigl{(}D(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),D\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)}\Bigr{)}$
$=H\Bigl{(}T\bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)}\Bigr{)}H\Bigl{(}D\bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)}\Bigr{)}.$
By repeated application of Lemma A.4, we have that
$\mathfrak{s}(T(T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}))=\mathfrak{s}(\mathbf{W}_{k+2})$
$=\mathfrak{s}(T(T(\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})),T(\mathbf{W}_{k+1},\mathbf{W}_{k}))),\quad\text{and}$
$\mathfrak{s}(T(D(T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k})),D(\mathbf{W}_{k+2},\mathbf{W}_{k+1})))$
$=\mathfrak{s}(\mathbf{W}_{k+1})=\mathfrak{s}(T\bigl{(}D(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),D\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)})).$
Applying Lemma 3.1, we conclude (3.2) and (3.3), as we wanted to prove. ∎
Proof of Lemma 3.1..
Immediately, the backwards direction is clear. Then, it remains to show that, under the assumption that $H(\mathbf{Y}_{1})H(\mathbf{Y}_{2})=H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}_{2}^{\prime})\implies\mathbf{Y}_{1}=\mathbf{Y}^{\prime}_{1},\mathbf{Y}_{2}=\mathbf{Y}_{2}^{\prime}$. Let us prove by backwards induction on $k=N,\ldots,1$ that $Y_{1,k}=Y^{\prime}_{1,k}$ and $Y_{2,k}=Y^{\prime}_{2,k}$.
As a base case, let’s show this for $k=N$. From Lemma 2.6, note that
$\left(H(\mathbf{Y}_{1})H(\mathbf{Y}_{2})\right)_{1,N}$ $=Z^{\mathbf{Y}}(2,N\mid 1,1)$
$=e^{Y_{2,N}}\left(Z^{\mathbf{Y}}(1,N\mid 1,1)+Z^{\mathbf{Y}}(2,N-1\mid 1,1)\right)$
$=e^{Y_{2,N}}\Biggl{(}\prod_{i=1}^{N}e^{Y_{1,i}}+\left(H(\mathbf{Y}_{1})H(\mathbf{Y}_{2})\right)_{1,N-1}\Biggr{)}.$ (3.6)
Similarly, we conclude that:
$\left(H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}^{\prime}_{2})\right)_{1,N}$ $=Z^{\mathbf{Y}^{\prime}}(2,N\mid 1,1)$
$=e^{Y^{\prime}_{2,N}}\left(Z^{\mathbf{Y}^{\prime}}(1,N\mid 1,1)+Z^{\mathbf{Y}^{\prime}}(2,N-1\mid 1,1)\right)$
$=e^{Y^{\prime}_{2,N}}\Biggl{(}\prod_{i=1}^{N}e^{Y^{\prime}_{1,i}}+\left(H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}^{\prime}_{2})\right)_{1,N-1}\Biggr{)}.$ (3.7)
By assumption, we have $H(\mathbf{Y}_{1})H(\mathbf{Y}_{2})=H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}_{2}^{\prime})$ and $\prod_{i=1}^{N}e^{Y_{1,i}}=e^{\mathfrak{s}(Y_{1})}=e^{\mathfrak{s}(Y^{\prime}_{1})}=\prod_{i=1}^{N}e^{Y^{\prime}_{1,i}}$. Equating (3.6) and (3.7) and canceling the nonzero expressions in parenthesis yields $Y_{2,N}=Y^{\prime}_{2,N}$. Next, we see that $Y^{\prime}_{1,N}=Y_{1,N}$, since:
$(H(\mathbf{Y}_{1})H(\mathbf{Y}_{2}))_{(N,N)}=Z^{\mathbf{Y}}(2,N\mid 1,N)=e^{Y_{1,N}+Y_{2,N}},$ (3.8)
and
$(H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}_{2}^{\prime}))_{(N,N)}=Z^{\mathbf{Y}^{\prime}}(2,N\mid 1,N)=e^{Y^{\prime}_{1,N}+Y^{\prime}_{2,N}}.$ (3.9)
To finish the proof, simply notice that the proof of the above only required $\prod_{i=1}^{N}e^{Y^{\prime}_{i}}=\prod_{i=1}^{N}e^{Y_{i}}$. Alternatively, if we have $\prod_{i=1}^{k}e^{Y_{1,i}}=\prod_{i=1}^{k}e^{Y_{2,i}}$ for some $k\in\llbracket 1,N\rrbracket$, then the same proof shows that $Y_{1,k}=Y_{1,k}^{\prime}$ and $Y_{2,k}=Y_{2,k}^{\prime}$. This follows by backwards induction, since
$\prod_{i=1}^{k}e^{Y_{1,i}}=\mathfrak{s}(\mathbf{Y}_{1})\Bigl{(}\prod_{i=k+1}^{N}e^{Y_{1,i}}\Bigr{)}^{-1}.\qed$
4. Burke property and permutation invariance
In this section, we prove Theorems 1.7 and 1.8. We first need two intermediate lemmas. In the first, we generalize the Burke property in [CGS24, Proposition 5.12] to an inhomogeneous environment.
Proposition 4.1 (Log-inverse-gamma Burke property).
Let $\lambda>0$, and let $\mathbf{a},\mathbf{b}\in\mathbb{R}^{\mathbb{Z}_{N}}$ be sequences such that $a_{i}+b_{j}>0$ for $i,j\in\mathbb{Z}_{N}$. Let $\mathbf{X}_{1},\mathbf{X}_{2}$ be independent sequences of independent random variables such that $X_{i,j}\sim(a_{i}+b_{j},\lambda)$ (see Definition 1.15). Then,
$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}(T(\mathbf{X}_{1},\mathbf{X}_{2}),D(\mathbf{X}_{1},\mathbf{X}_{2})).$
Proof.
Using Lemmas A.4-A.5 in the penultimate equality below, the joint density $\mathbb{P}(\mathbf{X}_{1}\in d\mathbf{x}_{1},\mathbf{X}_{2}\in d\mathbf{x}_{2})$ is proportional to
$\exp\Biggl{(}-\sum_{i\in\mathbb{Z}_{N}}\Bigl{(}(a_{1}+b_{i})x_{1,i}+(a_{2}+b_{i})x_{2,i}\Bigr{)}-\lambda\sum_{i\in\mathbb{Z}_{N}}(e^{-x_{1,i}}+e^{-x_{2,i}})\Biggr{)}$
$=\exp\Biggl{(}-a_{1}\sum_{i\in\mathbb{Z}_{N}}x_{1,i}-a_{2}\sum_{i\in\mathbb{Z}_{N}}x_{2,i}-\sum_{i\in\mathbb{Z}_{N}}b_{i}(x_{1,i}+x_{2,i})-\lambda\sum_{i\in\mathbb{Z}_{N}}(e^{-x_{1,i}}+e^{-x_{2,i+1}})\Biggr{)}$
$=\exp\Biggl{(}-a_{1}\sum_{i\in\mathbb{Z}_{N}}T_{i}(\mathbf{x}_{1},\mathbf{x}_{2})-a_{2}\sum_{i\in\mathbb{Z}_{N}}D_{i}(\mathbf{x}_{1},\mathbf{x}_{2})-\sum_{i\in\mathbb{Z}_{N}}b_{i}(T_{i}(\mathbf{x}_{1},\mathbf{x}_{2})+D_{i}(\mathbf{x}_{1},\mathbf{x}_{2}))$
$\qquad\qquad\qquad\qquad-\lambda\sum_{i\in\mathbb{Z}_{N}}(e^{-T_{i+1}(\mathbf{x}_{1},\mathbf{x}_{2})}+e^{-D_{i}(\mathbf{x}_{1},\mathbf{x}_{2})})\Biggr{)}$
$=\exp\Biggl{(}-\sum_{i\in\mathbb{Z}_{N}}\Bigl{(}(a_{1}+b_{i})T_{i}(\mathbf{x}_{1},\mathbf{x}_{2})+(a_{2}+b_{i})D_{i}(\mathbf{x}_{1},\mathbf{x}_{2})\Bigr{)}-\lambda\sum_{i\in\mathbb{Z}_{N}}(e^{-T_{i}(\mathbf{x}_{1},\mathbf{x}_{2})}+e^{-D_{i}(\mathbf{x}_{1},\mathbf{x}_{2})})\Biggr{)},$
Thus, the density is preserved under our transformation, so all we must show is that the determinant of the Jacobian matrix of the transformation has absolute value 1. This follows from Lemma A.7. ∎
We next prove a lemma that connects the condition stated in Theorem 1.7 to the condition in Theorem 1.4.
Lemma 4.2.
Let $\sigma:\mathbb{Z}\to\mathbb{Z}$ be a finite permutation, decomposed as a product of nearest-neighbor transpositions $(k_{m},k_{m}+1)(k_{m-1},k_{m-1}+1)\cdots(k_{1},k_{1}+1)$, where $k_{1},\ldots,k_{n}\in\mathbb{Z}$. Assume that this decomposition has the minimal number of terms. Then, if $x\in\mathbb{Z}$ satisfies $\sigma(\mathbb{Z}_{<x})=\mathbb{Z}_{<x}$, then $x\neq k_{\ell}+1$ for all $1\leq\ell\leq n$. Similarly, if $\sigma(\mathbb{Z}_{>x})=\mathbb{Z}_{>x}$, then $x\neq k_{\ell}$ for all $1\leq\ell\leq m$.
Proof.
We prove the first statement; the second has a symmetric proof. If $\sigma(\mathbb{Z}_{<x})=\mathbb{Z}_{<x}$, then $\sigma$ can be decomposed as $\sigma=\sigma_{1}\sigma_{2}$, where $\sigma_{1}$ permutes only elements of $\mathbb{Z}_{<x}$, and $\sigma_{2}$ permutes only elements of $\mathbb{Z}_{\geq x}$. By minimality of the decomposition $\sigma=(k_{m},k_{m}+1)(k_{m-1},k_{m-1}+1)\cdots(k_{1},k_{1}+1)$, we must have that
$\sigma_{1}=(k_{\ell_{r}},k_{\ell_{r}}+1)\cdots(k_{\ell_{1}},k_{\ell_{1}}+1),\quad\text{and}\quad\sigma_{2}=(k_{p_{n}},k_{p_{n}}+1)\cdots(k_{p_{1}},k_{p_{1}}+1)$
for disjoint subsequences $k_{\ell_{1}},\ldots,k_{\ell_{r}}$ and $k_{p_{1}},\ldots,k_{p_{n}}$ with $r+n=m$. Since $\sigma_{1}$ only permutes elements of $\mathbb{Z}_{<x}$, we have that $k_{\ell_{t}}+1<x$ for $1\leq t\leq r$. Since $\sigma_{2}$ only permutes elements of $\mathbb{Z}_{\geq x}$, we have $k_{p_{t}}+1>x$ for $1\leq t\leq n$. Hence, $x\neq k_{\ell}+1$ for $1\leq\ell\leq m$. ∎
Proof of Theorem 1.7..
For an arbitrary finite permutation $\sigma$, decompose it as a product of nearest-neighbor transpositions $\sigma=(k_{m},k_{m}+1)(k_{m-1},k_{m-1}+1)\cdots(k_{1},k_{1}+1)$. Take this decomposition to have the minimal number of terms. Then, we consider the action of the map $\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}$ on the weights $\mathbf{W}^{\mathbf{a},\mathbf{b}}$, as in Theorem 1.4. By Lemma 4.2, the condition (1.11) and the minimality of the decomposition ensures that the conditions of Theorem 1.4 are met, so
$Z^{\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}\mathbf{W}^{\mathbf{a},\mathbf{b}}}(V\mid U)=Z^{\mathbf{W}^{\mathbf{a},\mathbf{b}}}(V\mid U),\quad\text{for all }(U,V)\in\mathcal{U}^{\sigma}.$
Furthermore, by successively applying the Burke property in Proposition 4.1, we see that each operator $\mathcal{P}_{k}$ swaps the column parameters in columns $k$ and $k+1$, so we have that $\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}\mathbf{W}^{\mathbf{a},\mathbf{b}}\overset{d}{=}\mathbf{W}^{\sigma\mathbf{a},\mathbf{b}}$. This completes the proof. ∎
Proof of Theorem 1.8..
This is proved similarly as Theorem 1.7. Consider the independent weights in the finite grid $\llbracket-M,M\rrbracket\times\llbracket-M,M\rrbracket$, and extend these periodically. Consider first endpoint pairs $(U,V)$ that consist of points lying in the finite grid $\llbracket-M,M\rrbracket\times\llbracket-M,M\rrbracket$, and apply the partition function invariance in Theorem 1.4 and the Burke property in Proposition 4.1 (with Lemma 4.2 to connect the two conditions) to both rows and columns, with $N=2M+1$. The result follows by sending $M\to\infty$. ∎
Appendix A Inputs from previous work and technical inputs
A.1. Basic algebraic properties
In the present work, we use a different formulation of the maps $D$ and $T$ than those introduced in [CGS24]; there, the relevant mappings were defined as
$\mathfrak{T}_{i}(\mathbf{\mathbf{X}_{1}},\mathbf{\mathbf{X}_{2}})$ $=X_{1,i}+\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,[i,j]}-X_{1,[i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,[i+1,j]-X_{1,[i+1,j]}}}}\Biggr{)},\quad\text{and}$
$\mathfrak{D}_{i}(\mathbf{\mathbf{X}_{1}},\mathbf{\mathbf{X}_{2}})$ $=X_{2,i}+\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,(i,j]}-X_{1,(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,(i-1,j]}-X_{1,(i-1,j]}}}\Biggr{)}.$
In [CGS24], these maps were denoted as $T^{N,2}$ and $D^{N,2}$, respectively. It is important to note that this is equivalent to our definition of the maps defined in (1.2) up to a shift:
$D(\mathbf{X}_{1},\mathbf{X}_{2})=\mathfrak{D}(\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2}),\quad\text{and}\quad T(\mathbf{X}_{1},\mathbf{X}_{2})=\tau_{-1}\mathfrak{T}(\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2}).$ (A.1)
where, for $k\in\mathbb{Z}_{N}$, $\tau_{k}:\mathbb{R}^{\mathbb{Z}_{N}}\rightarrow\mathbb{R}^{\mathbb{Z}_{N}}$ is the shift map: $(\tau_{k}X)_{i}=X_{i+k}$ for all $i\in\mathbb{Z}_{N}$.
Lemma A.1.
As $\beta\to\infty$, $\frac{1}{\beta}\log\Bigl{(}\sum\limits_{i=0}^{k}\exp(\beta E_{i})\Bigr{)}$ converges to $\max\limits_{i}E_{i}$, uniformly as a function of the $E_{i}$ (but not as a function of $k$)
Proof.
First, observe that
$\frac{1}{\beta}\log\Bigl{(}\sum\limits_{i=0}^{k}\exp(\beta E_{i})\Bigr{)}-\max\limits_{i}E_{i}$ $=\frac{1}{\beta}\log\Bigl{(}\sum\limits_{i=0}^{k}\exp(\beta E_{i})\Bigr{)}-\frac{1}{\beta}\log\Bigl{(}\exp(\beta\max\limits_{i}E_{i})\Bigr{)}$
$=\frac{1}{\beta}\log\Biggl{(}\frac{\sum\limits_{i=0}^{k}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}.$
From here, we bound below and above.
$\frac{1}{\beta}\log\Biggl{(}\frac{\sum\limits_{i=0}^{k}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}$ $\leq\frac{1}{\beta}\log\Biggl{(}\frac{k\max\limits_{i}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}=\frac{1}{\beta}\log(k).$
Similarly,
$\frac{1}{\beta}\log\Biggl{(}\frac{\sum\limits_{i=0}^{k}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}$ $\geq\frac{1}{\beta}\log\Biggl{(}\frac{\max\limits_{i}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}=\frac{1}{\beta}\log(1)=0.$
This concludes our proof. ∎
Lemma A.2.
The maps $\mathcal{P}$ and $\widetilde{\mathcal{P}}$ defined in (1.2) and (1.13) are involutions on $\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$.
Proof.
By [CGS24, Corollary 5.9], the map $\mathcal{W}:\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}\to\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$ defined by
$\mathcal{W}(\mathbf{X}_{1},\mathbf{X}_{2})=\bigl{(}\mathfrak{D}(\mathbf{X}_{1},\mathbf{X}_{2}),\mathfrak{T}(\mathbf{X}_{1},\mathbf{X}_{2})\bigr{)},$
is an involution. Therefore, combined with (A.1), for $(\mathbf{W}_{1},\mathbf{W}_{2})\in\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$, we have
$\mathcal{P}(\mathcal{P}(\mathbf{W}_{1},\mathbf{W}_{2}))$ $=(T(D(\mathbf{W}_{2},\mathbf{W}_{1}),T(\mathbf{W}_{2},\mathbf{W}_{1})),D(D(\mathbf{W}_{2},\mathbf{W}_{1}),T(\mathbf{W}_{2},\mathbf{W}_{1})))$
$=(T(\mathfrak{D}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1}),\tau_{-1}\mathfrak{T}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1})),D(\mathfrak{D}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1}),\tau_{-1}\mathfrak{T}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1})))$
$=(\tau_{-1}\mathfrak{T}(\mathfrak{D}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1}),\tau_{1}\tau_{-1}\mathfrak{T}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1})),\mathfrak{D}(\mathfrak{D}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1}),\tau_{1}\tau_{-1}\mathfrak{T}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1})))$
$=(\tau_{-1}\mathfrak{T}(\mathfrak{D}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1}),\mathfrak{T}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1})),\mathfrak{D}(\mathfrak{D}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1}),\mathfrak{T}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1})))$
$=(\tau_{-1}\tau_{1}\mathbf{W}_{1},\mathbf{W}_{2})$
$=(\mathbf{W}_{1},\mathbf{W}_{2}).$
Thus, $\mathcal{P}$ is an involution. Passing to zero-temperature, we get that
$\widetilde{\mathcal{P}}(\widetilde{\mathcal{P}}(\mathbf{W}_{1},\mathbf{W}_{2}))$ $=\lim_{\beta\to\infty}\frac{1}{\beta}\mathcal{P}(\beta\widetilde{\mathcal{P}}(\mathbf{W}_{1},\mathbf{W}_{2}))$
$=\lim_{\beta\to\infty}\frac{1}{\beta}\mathcal{P}(\mathcal{P}(\beta\mathbf{W}_{1},\beta\mathbf{W}_{2}))$
$=\lim_{\beta\to\infty}\frac{1}{\beta}(\beta\mathbf{W}_{1},\beta\mathbf{W}_{2})$
$=(\mathbf{W}_{1},\mathbf{W}_{2}),$
where the penultimate equality follows from $\mathcal{P}$ being an involution and the rest follow from the uniform convergence of $\mathcal{P}$ via Lemma A.1. ∎
Lemma A.3.
$\widetilde{\mathcal{P}}|_{\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}\times\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}}$ is an involution on $\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}\times\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}$.
Proof.
Since $\mathcal{P}$ is an involution on $\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$, it suffices to show that $(\mathbf{X}_{1},\mathbf{X}_{2})\in\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}\times\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}$ implies $\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2})\in\mathbb{R}_{\geq 0}^{\mathbb{Z}_{N}}$ and $\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})\in\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}$. Fix any $i\in\mathbb{Z}_{N}$, and take $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}$. Set $Y_{i}=X_{2,i+1}-X_{1,i}$ for all $i\in\mathbb{Z}_{N}$. We observe that
$Y_{(i-1,j]}=\begin{cases}Y_{i}+Y_{(i,j]},&\text{ if }j\in\mathbb{Z}_{N}\setminus\{i-1\}\\ 0,&\text{ if }j=i-1\end{cases}$
It follows that:
$\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}$ $=0\vee\left(Y_{i}+\max_{j\in\mathbb{Z}_{N}\setminus\{i-1\}}Y_{(i,j]}\right)$
$\leq 0\vee\left(Y_{i}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}\right)$
$\leq 0\vee\left(X_{1,i}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}\right)$
$\leq X_{1,i}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}.$
Similarly, we have that $Y_{[i,j]}=\begin{cases}-Y_{i-1}+Y_{[i-1,j]},&\text{ if }j\in\mathbb{Z}_{N}\setminus\{i-1\}\\ Y_{[i-1,i-2]},&\text{ if }j=i-1\end{cases}$
Therefore, we know
$\max_{j\in\mathbb{Z}_{N}}Y_{[i,j]}$ $=Y_{[i-1,i-2]}\vee\left(-Y_{i-1}+\max_{j\in\mathbb{Z}_{N}\setminus\{i-1\}}Y_{[i-1,j]}\right)$
$\leq\max_{j\in\mathbb{Z}_{N}}Y_{[i-1,j]}\vee\left(X_{2,i-1}+\max_{j\in\mathbb{Z}_{N}}Y_{[i-1,j]}\right)$
$\leq X_{2,i-1}+\max_{j\in\mathbb{Z}_{N}}Y_{[i-1,j]}.$
By the definitions in (1.13), we see that $\widetilde{T}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})\geq 0$ and $\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})\geq 0$ for all $i\in\mathbb{Z}_{N}$. ∎
Lemma A.4.
Let $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}$. Then,
  1. (i)
    $\mathfrak{s}(T(\mathbf{X}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{1})$
  2. (ii)
    $\mathfrak{s}(D(\mathbf{X}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{2})$
  3. (iii)
    $\mathfrak{s}(\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{1})$
  4. (iv)
    $\mathfrak{s}(\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{2})$
Proof.
These all follow immediately from the definitions (1.2) and (1.13) since the sums are telescoping. ∎
Lemma A.5.
[CGS24, Lemma 5.10] Let $\mathbf{X}_{1},\mathbf{X_{2}}\in\mathbb{R}^{\mathbb{Z}_{N}}$. Then, for $i\in\mathbb{Z}_{N}$,
  1. (i)
    $D_{i}(\mathbf{X}_{1},\mathbf{X_{2}})+T_{i}(\mathbf{X}_{1},\mathbf{X_{2}})=X_{1,i}+X_{2,i}$
  2. (ii)
    $e^{-D_{i}(\mathbf{X}_{1},\mathbf{X_{2}})}+e^{-T_{i+1}(\mathbf{X}_{1},\mathbf{X_{2}})}=e^{-X_{1,i}}+e^{-X_{2,i+1}}$
  3. (iii)
    $\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X_{2}})+\widetilde{T}_{i}(\mathbf{X}_{1},\mathbf{X_{2}})=X_{1,i}+X_{2,i}$.
Proof.
(i)-(ii) were proved in shifted form in [CGS24]. (iii) is the zero-temperature limit of (i). ∎
Lemma A.6.
[CGS24, Lemma B.3] Let $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}$.
  1. (i)
    If $(\mathbf{U}_{1},\mathbf{U}_{2})=\mathcal{P}(\mathbf{X}_{1},\mathbf{X}_{2})$, then $\mathcal{P}(\tau_{1}\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2})=(\tau_{1}\mathbf{U}_{1},\tau_{1}\mathbf{U}_{2})$.
  2. (ii)
    If $(\mathbf{U}_{1},\mathbf{U}_{2})=\widetilde{\mathcal{P}}(\mathbf{X}_{1},\mathbf{X}_{2})$, then $\widetilde{\mathcal{P}}(\tau_{1}\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2})=(\tau_{1}\mathbf{U}_{1},\tau_{1}\mathbf{U}_{2})$.
Lemma A.7.
The determinant of the Jacobian of the transformation
$\mathcal{P}:(\mathbf{X}_{1},\mathbf{X}_{2})\to(T(\mathbf{X}_{2},\mathbf{X}_{1}),D(\mathbf{X}_{2},\mathbf{X}_{1}))$
has absolute value $1$.
Proof.
This is shown in [CGS24, Proposition 5.12]. The version of the periodic Pitman transform used there is a shifted version of the map we use, which amounts to swapping columns and rows of the Jacobian matrix, leaving the absolute value of the determinant unchanged. ∎
Lemma A.8.
For $\mathbf{X}_{1},\mathbf{X}_{2},\mathbf{X}_{3}\in\mathbb{R}^{\mathbb{Z}_{N}}$,
$D(\mathbf{X}_{1},D(\mathbf{X}_{2},\mathbf{X}_{3}))=D\bigl{(}D(\mathbf{X}_{1},\mathbf{X}_{2}),D(T(\mathbf{X}_{1},\mathbf{X}_{2}),\mathbf{X}_{3}).$
Proof.
We know from [CGS24, Proposition 5.13] that
$\mathfrak{D}(\mathbf{X}_{1},\mathfrak{D}(\mathbf{X}_{2},\mathbf{X}_{3}))=\mathfrak{D}\bigl{(}\mathfrak{D}(\mathbf{X}_{1},\mathbf{X}_{2}),\mathfrak{D}(\mathfrak{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\mathbf{X}_{3}\bigr{)}.$ (A.2)
By (A.1) and Lemma A.6, we have
$\tau_{1}\mathfrak{D}(\mathbf{X}_{1},\mathbf{X}_{2})=\mathfrak{D}(\tau_{1}\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2}).$ (A.3)
Then, we have the following:
$\quad D\bigl{(}D(\mathbf{X}_{1},\mathbf{X}_{2}),D(T(\mathbf{X}_{1},\mathbf{X}_{2}),\mathbf{X}_{3})\bigr{)}$
[…]
as desired. ∎
Appendix B Proofs of the zero-temperature analogues in Theorems 1.10-1.12
Because polymer models have a strong connection to last-passage percolation (LPP), our algebraic results will naturally generalize to zero-temperature equivalents, through limiting arguments. In these limits, the $(+,\times)$ algebra is swapped for the $(\max,+)$ algebra, and our maps reflect this change. As introduced in [CGS24], the ‘correct’ map to look at here is $\widetilde{D}$, the uniform limit as $\beta$ goes to infinity of $\frac{1}{\beta}D(\beta\mathbf{X}_{1},\beta\mathbf{X}_{2})$, as well as $\widetilde{T}$, which is reached through an identical limit. In this section, we use our algebraic results about the behavior of the polymer model under the periodic Pitman transform (as well as the aforementioned limiting arguments) to prove analogous facts about LPP under the zero-temperature periodic Pitman transform. The following intermediate result is needed.
Lemma B.1.
For any $k\in\mathbb{N}$ and any pair of $k$-tuples of ordered pairs $U=\bigl{(}(a_{i},b_{i})\bigr{)}_{i\in\llbracket 1,k\rrbracket}$, $V=\bigl{(}(c_{i},d_{i})\bigr{)}_{i\in\llbracket 1,k\rrbracket}$, as $\beta\to\infty$, $\frac{1}{\beta}\log\bigl{(}Z^{\beta,\mathbf{W}}(V\mid U)\bigr{)}$ converges uniformly (as a function of $\mathbf{W}$) to $G^{\mathbf{W}}(V\mid U)$.
Proof.
This follows as a direct consequence of Lemma A.1 since there are finitely many multi-paths between any two sets of points. ∎
Note that in the special case where $U,V$ are points, this shows that the single-path polymer partition function converges to single-path LPP. Also, because the convergence to LPP is uniform on the weights, we know that the zero-temperature limit is well-behaved up to weights which depend on $\beta$. If the weights $\mathbf{W}_{\beta}$ converge to $\mathbf{\widetilde{\mathcal{W}}}$, then $\frac{1}{\beta}\log(Z^{\beta,\mathbf{W}_{\beta}}(V\mid U))$ will converge to $G^{\mathbf{\widetilde{\mathcal{W}}}}(V\mid U)$. With this in mind, we can now prove Theorem 1.10.
Proof of Theorem 1.10..
As in Theorem 1.4, we take $k_{1},k_{2},\ldots,k_{m}\in\mathbb{Z},U=\bigl{(}a_{i},b_{i})_{i\in\llbracket 1,m\rrbracket}\bigr{)}$, and $V=\bigl{(}(c_{i},d_{i})_{i\in\llbracket 1,m\rrbracket}\bigr{)}$ satisfying $(U,V)\in\Psi$ and, for all $i,j\in\llbracket 1,n\rrbracket$, $a_{i}\neq k_{j}+1$ and $c_{i}\neq k_{j}$. To prove Theorem 1.10, we must show that $G^{\mathbf{W}}(V\mid U)=G^{\widetilde{\mathcal{P}}_{k_{n}}\ldots\widetilde{\mathcal{P}}_{k_{1}}\mathbf{W}}(V\mid U)$, where the operators $\mathcal{P}_{k}$ are defined in 1.9, with $T,D$ replaced by $\widetilde{T},\widetilde{D}$. Now, for $\beta>0$, let $\mathcal{P}_{k}^{\beta}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ be operators defined by their action:
$\quad\,\mathcal{P}_{k}^{\beta}(\ldots,\mathbf{W}_{k-1},\mathbf{W}_{k},\mathbf{W}_{k+1},\mathbf{W}_{k+2},\ldots)$ (B.1)
$=(\ldots,\mathbf{W}_{k-1},\frac{1}{\beta}T(\beta\mathbf{W}_{k+1},\beta\mathbf{W}_{k}),\frac{1}{\beta}D(\beta\mathbf{W}_{k+1},\beta\mathbf{W}_{k}),\mathbf{W}_{k+2},\ldots).$
By Theorem 1.4, we know that $Z^{\mathbf{W}}(V\mid U)=Z^{\mathcal{P}_{k_{m}}\ldots\mathcal{P}_{k_{1}}\mathbf{W}}(V\mid U)$. Then, by definition of the polymer model with inverse temperature $\beta$ (1.8), for all $\beta>0$,
$Z^{\beta,\mathbf{W}}(V\mid U)=Z^{\beta\mathbf{W}}(V\mid U)=Z^{\mathcal{P}_{k_{1}}(\beta\mathbf{W})}(V\mid U)=Z^{\beta(\frac{1}{\beta}\mathcal{P}_{k_{1}}(\beta\mathbf{W}))}(V\mid U)=Z^{\beta,\mathcal{P}_{k_{1}}^{\beta}(\mathbf{W})}(V\mid U).$ (B.2)
By induction, we have $Z^{\beta,\mathbf{W}}(V\mid U)=Z^{\beta,\mathcal{P}_{k_{n}}^{\beta}\ldots\mathcal{P}_{k_{1}}^{\beta}\mathbf{W}}(V\mid U)$.
By Lemma B, we know that $\frac{1}{\beta}\log\bigl{(}Z^{\beta,\mathbf{W}}(V\mid U)\bigr{)}$ converges to $G^{\mathbf{W}}(V\mid U)$ uniformly on the weights, so $G^{\mathbf{W}}(V\mid U)$ is preserved under the limiting operator $\lim\limits_{\beta\to\infty}\mathcal{P}_{k_{n}}^{\beta}\ldots\mathcal{P}_{k_{1}}^{\beta}$. We claim that each $\mathcal{P}_{k}^{\beta}$ converges to $\widetilde{\mathcal{P}}_{k}$ uniformly as a function of the weights, which implies by induction that
$G^{\mathbf{W}}(V\mid U)=G^{\lim\limits_{\beta\to\infty}\mathcal{P}_{k_{m}}^{\beta}\ldots\mathcal{P}_{k_{1}}^{\beta}\mathbf{W}}(V\mid U)=G^{\widetilde{\mathcal{P}}_{k_{m}}\ldots\widetilde{\mathcal{P}}_{k_{1}}\mathbf{W}}(V\mid U),$ (B.3)
proving the theorem. The necessary uniform convergence is shown component-wise:
$\quad\,\frac{1}{\beta}D_{i}(\beta\mathbf{X}_{1},\beta\mathbf{X}_{2})-\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})$
$=X_{2,i+1}+\frac{1}{\beta}\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i-1,j]}}}\Biggr{)}-(X_{2,i+1}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]})$
$=\Bigl{(}\frac{1}{\beta}\log\bigl{(}\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i,j]}}\bigr{)}-\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}\Bigr{)}-\Bigl{(}\frac{1}{\beta}\log\bigl{(}\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i-1,j]}}\bigr{)}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}\Bigr{)}.$
By Lemma A.1, we know that $\frac{1}{\beta}\log(N)\leq\frac{1}{\beta}D_{i}(\beta\mathbf{X}_{1},\beta\mathbf{X}_{2})-\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})\leq\frac{1}{\beta}\log(N)$, proving uniform convergence. An identical argument holds for $T_{i}-\widetilde{T}_{i}$. ∎
We next show the braid relations pass to zero-temperature.
Proof of Theorem 1.11..
Defining $\mathcal{P}_{k}^{\beta}$ as in (B.1), we see that $\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}=\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}$, as an immediate consequence of Theorem 1.6. As noted in the proof of Theorem 1.10, $\mathcal{P}_{k}^{\beta}$ converges to $\widetilde{\mathcal{P}}_{k}$ uniformly. This implies the braid relation for $\widetilde{\mathcal{P}}_{k}$, since we have
$\widetilde{\mathcal{P}}_{k}\widetilde{\mathcal{P}}_{k+1}\widetilde{\mathcal{P}}_{k}=\lim\limits_{\beta\to\infty}\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}=\lim\limits_{\beta\to\infty}\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}=\widetilde{\mathcal{P}}_{k+1}\widetilde{\mathcal{P}}_{k}\widetilde{\mathcal{P}}_{k+1}.\qed$
To prove Theorem 1.12, we first need analogous Burke properties in the geometric and exponential cases.
Proposition B.2 (Geometric Burke property).
Take $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{Z}_{\geq 0}^{\mathbb{Z}_{N}}$ to be independent sequences of independent random variables such that $X_{1,i}\sim\text{Geom}(p_{1}q_{i})$ and $X_{2,i}\sim\text{Geom}(p_{2}q_{i})$ for all $i\in\mathbb{Z}_{N}$. Then,
$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}(\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})).$
Proof.
Letting $\mathbf{k}_{1},\mathbf{k}_{2}\in\mathbb{Z}_{\geq 0}^{N}$ where $k_{m,i}$ is the $i$th component of $\mathbf{k}_{m}$, we have that
$\mathbb{P}(\mathbf{X}_{1}=\mathbf{k}_{1},\mathbf{X}_{2}=\mathbf{k}_{2})$ $=\mathbb{P}\left(X_{1,1}=k_{1,1}\right)\cdots\mathbb{P}\left(X_{1,N}=k_{1,N}\right)\mathbb{P}\left(X_{2,1}=k_{2,1}\right)\cdots\mathbb{P}\left(X_{2,N}=k_{2,N}\right)$
$=\left(\prod\limits_{i=1}^{N}(1-p_{1}q_{i})(1-p_{2}q_{i})\right)p_{1}^{\mathfrak{s}(\mathbf{k}_{1})}p_{2}^{\mathfrak{s}(\mathbf{k}_{2})}\prod_{i=1}^{N}q_{i}^{k_{1,i}+k_{2,i}}$
$=\left(\prod\limits_{i=1}^{N}(1-p_{1}q_{i})(1-p_{2}q_{i})\right)p_{1}^{\mathfrak{s}(\widetilde{D}(\mathbf{k}_{2},\mathbf{k}_{1}))}p_{2}^{\mathfrak{s}(\widetilde{T}(\mathbf{k}_{2},\mathbf{k}_{1}))}\prod_{i=1}^{N}q_{i}^{\widetilde{D}_{i}(\mathbf{k}_{2},\mathbf{k}_{1})+\widetilde{T}_{i}(\mathbf{k}_{2},\mathbf{k}_{1})}$
$=\prod\limits_{i=1}^{N}(1-p_{1}q_{i})(p_{1}q_{i})^{\widetilde{D}_{i}(\mathbf{k}_{2},\mathbf{k}_{1})}\prod\limits_{i=1}^{N}(1-p_{2}q_{i})(p_{2}q_{i})^{\widetilde{T}_{i}(\mathbf{k}_{2},\mathbf{k}_{1})}$
$=\mathbb{P}(\mathbf{X}_{1}=\widetilde{D}(\mathbf{k}_{2},\mathbf{k}_{1}),\mathbf{X}_{2}=\widetilde{T}(\mathbf{k}_{2},\mathbf{k}_{1}))$
$=\mathbb{P}(\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})=\mathbf{k}_{2},\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2})=\mathbf{k}_{1})$
where the first equality follows from the independence of $\mathbf{X}_{1},\mathbf{X}_{2}$ and $X_{m,i},X_{m,j}$ for $m\in\{1,2\}$ and $i,j\in\mathbb{Z}_{N}$ where $i\neq j$; the second equality follows from our definition of the geometric distribution; the third equality follows from Lemmas A.4(iii)-(iv) and A.5(i); and the final equality follows from the bijection in Lemma A.3. ∎
Proposition B.3 (Exponential Burke property).
Let $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{Z}_{\geq 0}^{\mathbb{Z}_{N}}$ be independent sequences of independent random variables such that $X_{1,i}\sim\text{Exp}(a_{1}+b_{i})$ and $X_{2,i}\sim\text{Exp}(a_{2}+b_{i})$ for all $i\in\mathbb{Z}_{N}$ where $a_{m}+b_{i}>0$ for $m\in\{1,2\}$ and $i\in\mathbb{Z}_{N}$. Then,
$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}(\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})).$
Proof.
For $m\in\{1,2\}$ and $i\in\mathbb{Z}_{N}$, define
$X_{m,i}^{\prime}\sim\text{Geom}\Biggl{(}\left(1-\frac{a_{m}}{n}\right)\left(1-\frac{b_{i}}{n}\right)\Biggr{)}.$
Then, since $\left(1-\frac{a_{m}}{n}\right)\left(1-\frac{b_{i}}{n}\right)=1-\frac{a_{m}+b_{i}}{n}+O(n^{-2})$, we have that
$\lim\limits_{n\to\infty}\frac{X_{m,i}^{\prime}}{n}\overset{d}{=}X_{m,i}\sim\text{Exp}(a_{m}+b_{i}),\quad\text{for }m\in\{1,2\},\text{ and }i\in\mathbb{Z}_{N}$
Observe from the definitions of $\widetilde{T}$ and $\widetilde{D}$ (1.13) that
$\frac{1}{n}\widetilde{T}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})=\widetilde{T}\Bigl{(}\frac{\mathbf{X}_{1}^{\prime}}{n},\frac{\mathbf{X}_{2}^{\prime}}{n}\Bigr{)},\quad\text{and}\quad\frac{1}{n}\widetilde{D}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})=\widetilde{D}\Bigl{(}\frac{\mathbf{X}_{1}^{\prime}}{n},\frac{\mathbf{X}_{2}^{\prime}}{n}\Bigr{)}.$
By Proposition B.2, we know that $(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})\overset{d}{=}(\widetilde{T}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}),\widetilde{D}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}))$. Since $\max$, addition, and subtraction are continuous, $\widetilde{D}$ and $\widetilde{T}$ are continuous maps. We then have that
$\Bigl{(}\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})\Bigr{)}\overset{d}{=}\lim\limits_{n\to\infty}\frac{1}{n}\Bigl{(}\widetilde{T}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}),\widetilde{D}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})\Bigr{)}\overset{d}{=}\lim\limits_{n\to\infty}\frac{1}{n}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})\overset{d}{=}(\mathbf{X}_{1},\mathbf{X}_{2}).\qed$
Proof of Theorem 1.12.
The proof is the same as the proof of Theorems 1.7 and 1.8: replace the use of Theorem 1.4 with Theorem 1.10 and replace the Burke property in Propositions 4.1 with the Burke property in Propositions B.2 and B.3. ∎